Area of a sail. The area in square meters for a triangular
sail is given by A(x) = x2 + 5x + 6.
a) Find A(5).
b) If the height of the sail is x + 3 meters, then what is
the length of the base of the sail?

Answer by RAY100(1637)

A(x)= x^2 +5x +6
.
a(5) =(5)^2 +5(5) +6= 25 +25 +6 = 56
.
A = 1/2 * b*h,,,,with h=(x+3)=(5)+3=8
.
A = 1/2 *8 *b = 56
.
4b=56
.
b=14,,,in general b=x+y, at x=5, b=14=5+y,,y=9,,,therefore b=(x+9)
.
check, A=1/2 *(x+3) *(x+9) = 1/2 *8*14= 56,,,ok